Simplify 3x (4x – 5) + 3 and find its values for (i) x = 3 (ii) x =1/2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3x (4x – 5) + 3
=3x(4x)–3x(5) +3
=12x^2- 15x + 3

(i) putting a= 3 in the equationwe get
12x^2- 15x + 3 =12(3^2)–15 (3) +3
= 108–45 + 3 = 66

(ii)putting a=1/2 in the equation we get
12x^2- 15x + 3 = 12 (1/2)^2 – 15 (1/2) + 3
=12 (1/4) - 15/2 +3
= 3 - 15/2 + 3
= 6- 15/2
= (12- 15 ) /2
= -3/2

Thanks